Molly has a jar with nickels, dimes and quarters in it. The ratio of the number of quarters to the number of dimes to the number of nickels in the jar is 9:3:1. The total value of all the coins in the Jar is $18.20. How many coins does molly have in her jar?
Answer by ikleyn(52798)

.
Molly has a jar with nickels, dimes and quarters in it.
The ratio of the number of quarters to the number of dimes to the number of nickels
in the jar is 9:3:1. The total value of all the coins in the Jar is $18.20.
How many coins does molly have in her jar?
~~~~~~~~~~~~~~~~~

There are x nickels, 3x dimes, and 9x quarters in the jar,
where x is the common measure (now unknown value).


The total money equation is

    5x + (3x)*10 + (9x)*25 = 1820  cents.


Simplify and find x

    5x + 30x + 225x = 1820

         260x = 1820

            x = 1820/260 = 7.


ANSWER.  7 nickels, 3*7 = 21 dimes and 63 quarters.


CHECK.  7*5 + 21*10 + 63*25 = 1820 cents, in total.    ! correct !

Solved.

================

Instead of writing equations, you could solve the problem mentally, using the grouping method.